OpenStudy (anonymous):

Help? Find all solutions to the equation. 7 sin^2x - 14 sin x + 2 = -5

OpenStudy (mertsj):

1. Add 5 to both sides 2. Divide both sides by 7 3. Factor the left side 4. Set each factor equal to 0 5. Solve each equation.

OpenStudy (anonymous):

Okay, so I was on the right track. I did all of that, up to #3. How do I factor sin^2x-2sin x?

OpenStudy (mertsj):

That's not what you get when you divide both sides by 7. What is 7 divided by 7?

OpenStudy (anonymous):

Oops. I'm an idiot.

OpenStudy (anonymous):

7 sin^2x - 14 sin x + 2 = -5 7 sin^2x - 14 sin x + 7 = 0 sin^2x - 2 sin x + 1 = 0 go from here.

OpenStudy (anonymous):

(sin x -1)^2

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

So sinx = 1

OpenStudy (anonymous):

So does that mean 1 is the only solution?

OpenStudy (mertsj):

The number of solutions is infinite unless you have a restricted domain.

OpenStudy (mertsj):

x=90 +360n where n is an integer.

OpenStudy (anonymous):

Okay, so infinite. Thanks. You made this so much more simple than my online class does. :)

OpenStudy (mertsj):

yw